OpenStudy (anonymous):

Paul's Plumbing charges a flat fee of $90 plus $40/h for a plumbing job. Perfect Pipes charges a flat fee of $60 plus $55/h for a plumbing job. A. Write expressions to represent the total costs of a plumbing job for each company. Let h represent the time, in hours, of a plumbing job. Be sure to label the expression for each company! B. Is the cost of a 3-hour-long plumbing job the same at both companies. Write and use an equation to find out. Show all work!

OpenStudy (anonymous):

(a) Paul's Plumbing: c = 90 + 40h Perfect pipes : c = 60 + 55h (b) h = 3

OpenStudy (anonymous):

am i doing this right ? @Love_Ranaa

OpenStudy (love_ranaa):

Oh I just did the same exact thing a few days ago.

OpenStudy (anonymous):

Oh haha really, i'm kylies friend im in k12 so im using her account (:

OpenStudy (love_ranaa):

I can show you what I did. And I'm in k12 too. 9th grade.

OpenStudy (anonymous):

Same here, i go to Georgia Cyber Academy

OpenStudy (love_ranaa):

lol same.

OpenStudy (anonymous):

its k12 too but yeah if you could show me please, and omg really, you should message me !

OpenStudy (love_ranaa):

Was it 1.06 or 1.09?

OpenStudy (anonymous):

Lesson 1.10

OpenStudy (love_ranaa):

It wont let me see it again but what I did was a) paul's plumbing: c= 90 x 40h perfect pipes: c= 60 x 50h

OpenStudy (love_ranaa):

sorry i meant + not x ^

OpenStudy (anonymous):

On both of them right?

OpenStudy (love_ranaa):

b) 90+40h h= 3 40 x 3 = 120 + 90= $210 perfect pipes: 60+50h h=3 50x3= 150+60=$210

OpenStudy (love_ranaa):

Yeah on both of them I mean + . So the cost is the same. I know I do a little extra work though :p

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (love_ranaa):

No problem

OpenStudy (anonymous):

Do you have a instagram?